Water flows through a fire hose of diameter 6.35 cm at a rate of 0.0115 m3/s. The fire hose ends in a nozzle of inner diameter 2.20 cm. What is the speed with which the water exits the nozzle?
? m/s
Answer #1

here,

the volume flow rate , V = 0.0115 m^3 /s

the inner diameter of nozzle , d = 2.2 cm = 0.022 m

radius , r = d/2 = 0.011 m

let the speed with which the water exits be v

volume flow rate , V = (pi * r^2) * v

0.0115 = pi * 0.011^2 * v

solving for v

v = 30.27 m/s

the speed with which the water exits is 30.27 m/s